OpenStudy (anonymous):

how can i solve tis math problem 3 1/3 * 2 3/5 simplify or reduce

Parth (parthkohli):

=> 10/3 * 13/5 => 130/15 => 26/3 => 8 2/3

OpenStudy (anonymous):

could you show me step by step please

Parth (parthkohli):

I have! 3 1/3 = 10/3 and 2 3/5 = 13/5 Hope this helps:)

OpenStudy (anonymous):

@abby12 do you know how to multiply fractions?

Parth (parthkohli):

@abby12 If not, here's it. \[\LARGE {a \over b} * {c \over d} = {ac \over cd}\]